OpenStudy (anonymous):

I did this question and got p=1 and -2 but i want a second opinion. Please show all working. The function f(x)=x^3 - p^2x^2 + 2x - p has remainder -5 when it is divided by x + 1. Find the possible values of the constant p.

OpenStudy (anonymous):

Given that The function f(x)=x^3 - p^2x^2 + 2x - p has remainder -5 when it is divided by x + 1. So we have f(-1) = -5 now plugging -1 in place of x in f(x) we get the quadratic plotynomial in p that is \[-1-p ^{2}-2-p=-5\] simplifying we get \[p ^{2}+p-2=0\] that gives us p = 1 and -2 Keep It up Crimson!!

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

Its My pleasure!

OpenStudy (anonymous):

yeah thts wat i got